OpenStudy (anonymous):

Which of the following is not a step used when graphing the equation y = −x + 8 using the slope-intercept method? Explain the correct order of steps needed to graph this equation, using the slope-intercept method.

OpenStudy (anonymous):

Correct order of steps: \[y=-x+8\] since \[y=mx+b\] is the slope-intercept form where m is the slope and b is the y-intercept: m=-1, b=8. plot (0,8) or 8 in the y-axis. next, use the slope to get another point needed to define the line. the point needed is (0+1,8-1)=(1,7), since -1=-1/1 (-1 is y, 1 is x). then draw the resulting line from tracing a line through the two plotted points :))

OpenStudy (anonymous):

your slope is -1, which is equal to delta y/delta x = -1/1